OpenStudy (anonymous):

@Nurali @nincompoop he population of the world was 5.7 billion in 1995, and the observed relative growth rate was 2% per year. By what years will the population double and triple?

OpenStudy (anonymous):

@Nurali u know how to do it

OpenStudy (nurali):

A = 5.7 • e^(0.02t) 2 • 5.7 = 5.7 • e^(0.02t) 2 = e^(0.02t) ln 2 = 0.02t 34.65735903... = t It wll double during the year 2029. (1995 + 34) 3 • 5.7 = 5.7 • e^(0.02t) 3 = e^(0.02t) ln 3 = 0.02 54.93061443... = t It will triple during the year 2049. (1995 + 54)

OpenStudy (anonymous):

@Nurali what does the 5.7 come in

OpenStudy (nurali):

in question 5.7 billion

OpenStudy (anonymous):

whats the equation again whats with the question marks?

OpenStudy (anonymous):

@Nurali u get what im saying

OpenStudy (anonymous):

@Nurali hello

OpenStudy (nurali):

Af=Aoe^(rt) 2=1e^(0.02t) 2=e^(0.02t) ln(2)/0.02=t t=34.65, so add 1995 and get 2029.65 ln(3)/.02= 54.93, add 1995 and get 2049.93